OpenStudy (anonymous):

Jamie has 300 ft of fencing to enclose a rectangular pasture. The pasture's length is to be 10 ft less than 3 times the width. Find the width of the garden.

OpenStudy (anonymous):

Jamie has 300 ft of fencing to enclose a rectangular pasture. The pasture's length is to be 10 ft less than 3 times the width. Find the width of the garden.

OpenStudy (hockeychick23):

ok so you have 300 in total, here is your equation you're going to use: L= 3w-10 So you have: w+w+3w-10+3w-10=300 So now you have: 8w-20= 300 Add 20 to both sides: 8w-20+20=300+20 Now you have: 8w=320 Divide both sides by 8: 8w/8=320/8 w=40 Now fill in w to the length equation L= 3(40)-10 L= 110 So your width is 40 and your length is 110

OpenStudy (arabpride):

Great! Please close this post if you're done. Thanks! :D